OpenStudy (anonymous):

Janine has job offers at two companies. One company offers a starting salary of $28,000 with the raise of $3000 each year. The other company offers a starting price of $36,000 with a raise if $2000 each year. a. After how many years would Janine salary be the same with both companies? b. What would that salary be?

OpenStudy (anonymous):

28,000+3,000x=36,000+2,000x 3,000x-2,000x=36,000-28,000 1,000x=8,000x=8 years the salaries will be the same. Proof: 28,000+3,000*8=36,000+2,000*8 28,000+24,000=36,000+16,000 $52,000=$52,000 The salary from both companies after 8 years.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Np sis :)